OpenStudy (anonymous):

Note: My school has a study guide, yet when referring to the text, it does not say anything about this... The difference between a relation and a function is _______________?

OpenStudy (anonymous):

I am in algebra by the way

OpenStudy (anonymous):

a function is a special type of relation..

OpenStudy (anonymous):

Thanks, I didnt see that wiki answer.

OpenStudy (anonymous):

what does your text say what is a function?

OpenStudy (anonymous):

For as long as you have taken math courses, you have worked with equations similar to the ones shown below. 1 + 2=3 x + 7 = 10 4x + 7 = y x2 – 7x + 6 = y Most of the equations you have worked with up to this point have also been functions. But, what's the difference between an equation and a function? Take a look at the linear equation y = 2x + 3. This is a linear equation because the degree of the equation is 1. The graph of this equation can be found by plotting the y-intercept at 3 and counting the slope 2 (rise over run). Or, you could choose x-coordinates to substitute into the original equation, and then find the corresponding y-coordinates. Let's take a look at this second method.

OpenStudy (anonymous):

Notice that each member of the domain (x-coordinate) corresponds to a unique member of the range (y-coordinate). There isn't a single x-coordinate that is associated with two different y-coordinates. If you think of the x-coordinates as being the input values and the y-coordinates as the output values, you could say that every input has only one distinct output. This is the definition of a function.

OpenStudy (anonymous):

that last post, second sentence... that's what a function is.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw...